OpenStudy (anonymous):

what is 2nd derivative of x+x^-1

OpenStudy (anonymous):

f(x) = x+x^-1 f'(x) = 1 -x^-2 f''(x) = 2x^-3

OpenStudy (anonymous):

Basically just take the first derivative of the original equation, then you take the derivative of the derivative , which is the second derivative.

OpenStudy (anonymous):

YIANO, LOOK AT MY PROBLEM AND HELP ME PLZ